OpenStudy (anonymous):

what is the x-intercept of 3x + y = 7

OpenStudy (raden):

take y = 0 then solve for x

OpenStudy (anonymous):

inorder to find x make y=0

OpenStudy (raden):

3x + y = 7 3x + 0 = 7 3x = 7 x = .... ?

OpenStudy (anonymous):

okay let me work that out

OpenStudy (anonymous):

would x= 7/3 or would it equal 2.3?

OpenStudy (raden):

completely, it is (7/3 , 0)

OpenStudy (anonymous):

okay! thank you so much!

OpenStudy (raden):

you're welcome :)